OpenStudy (anonymous):

What is the solution to the equation? -18+4+x=9-19

OpenStudy (anonymous):

-18+4+x=9-19

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

ok

pooja195 (pooja195):

−18+4+x=9−19 −18+4+x=9+−19 (x)+(−18+4)=(9+−19)(Combine Like Terms) x+−14=−10 x−14=−10 Add 14 to both sides to get answers

OpenStudy (anonymous):

@pooja195 i have a hard time with equations

pooja195 (pooja195):

-10+14=?

OpenStudy (anonymous):

both sides of what?

OpenStudy (anonymous):

in order to cancel the -14 you need to do this: x-14+14=-10+14 which then equals X=4

OpenStudy (anonymous):

@pooja195 when you have time is it ok if you help me understnad equations?

pooja195 (pooja195):

@MariahDimond :) feel free to message me with any questions ^_^